# LDIR: Low-Dimensional Dense and Interpretable Text Embeddings with Relative Representations

### v1

@@ 5 Analyses and Discussions
In our main experiments, we remove the long texts in corpus and use FPS to extract the anchor texts. We analysis the impact of sampling methods (Section 5.1) and the length of anchor texts (Section 5.2) below. Then we discuss how to obtain a more interpretable dense embeddings with the calculation of fine-grained relatedness (Section 5.3) and show a case study (Section 5.5) in the final.
@@ 5.4 Surface-based and Vector-based Metrics
We further tried surface-based metrics as well as other various vector-based metrics for calculating relatedness, and the results for 500 dimensional embeddings are shown in Table 8. We found that surface-based distances cannot effectively represent semantic relationships, where the results are very poor. On the other hand, other vector-based distance (even when applied to LDIR after binarization) consistently show close results, despite differences in distances and cosine distances in downstream tasks.

### v2

@@ 5 Analyses and Discussions
In our main experiments, we remove the long texts in corpus and use FPS to extract the anchor texts. We analysis the impact of sampling methods (Section 5.1) and the length of anchor texts (Section 5.2) below. Then we discuss how to obtain a more interpretable dense embeddings with the calculation of fine-grained relatedness (Section 5.3), compare different metrics on relatedness (Section 5.4). Finally, we show a case study (Section 5.5) and discuss the usage of relative representation based embeddings (Section 5.6).
@@ 5.4 Surface-based and Vector-based Metrics
We further try surface-based metrics as well as other various vector-based metrics for calculating the relatedness, and the results for 500 dimensional embeddings are shown in Table 8. We find that surface-based distances cannot effectively represent semantic relationships, where the results are very poor. On the other hand, other vector-based distance (even when applied to LDIR after binarization) consistently show close results, despite differences in distances calculation against the cosine distances in downstream tasks.